# Efficient Adaptable High-throughput Streaming Aggregation Engines

### v3

@@ VI Discussion and Future work
The selection of operators in the aggregation pipeline can easily incorporate other associative operators such as multiplication, bitwise exclusive-OR etc. This can be achieved by including the corresponding logic inside the $n$ and $n^{\prime}$ entities. The average operation is not explored beyond simulation to avoid shifting the focus to floating point division optimisation. This is without loss of generality, only requiring $P$ dividers, usually consuming a few hundred LUT/FFs according to the use case, and without a perceptible timing overhead [36].
Similarly, other types of averaging are possible by combining simultaneous associative operators together at the $n^{\prime}$ stage. As an example, combining multiplication and count with a root circuit, can provide the geometric mean. Another instance is with the selection mechanism inside the SWAG engine, which can be generalised to support the percentile operator to select arbitrary percentiles based on the pre-calculated cardinality. Other selection queries could use more specialised criteria. Future applications could exploit the additional information in the engine, such as the order of the sorted data, which has already been used in the distinct count algorithm (step $(c)$ in section II-B) to essentially make it an associative operation.
@@
In the comparison of section V-B, a unique aspect for the SWAG engine with groups is the single window, while still providing per group statistics. Per-group window SWAG, although supported, might be less trivial to efficiently exploit the presented architecture, as the sorting advantage might diminish when “recycling” more data on each window advance. Thus, it could be seen as an approximation for when the task involves a higher number of groups and there is a need for multiple windows. In practice, keeping live statistics for all groups may not be necessary [32], and a simple filtering step to isolate the group IDs would still dedicate the whole window to the desired group.
Yet, the notion of having a common window for grouped data is widely useful for other SWAG tasks. One such case is for supporting more complex SWAG queries without groups, where the group field can carry the value instead, for operating in the context of the values. This includes the mincount or maxcount operator (i.e. counting the instances of the minimum or maximum value within the window) [37], which can be implemented as a group-by query where the count operator is only activated on the first or last group, corresponding the minimum or maximum respectively.
@@
The presented engines and base pipeline can be simplified further according to the desired operator selection. If only associative operators (e.g. sum) are needed, then the requirement for the sorted input is only with respect to the group. For SWAG, this essentially means that the comparison width of the small-scale sorter can be constrained to that of the group to minimise its complexity. For non-SWAG, this would also be reflected in the accompanying sorter, or in software equivalents that provide the sorted stream. If selection operators such as the median are not present, then the cardinality logic (section III-D) is not needed. Hence, the disjoint prefix-scan steps can also be merged back. In this scenario, alternative sorters would also suffice for SWAG. When this holds and the input does not carry group information, then a sorter is not needed at all.
As demonstrated in the study of section IV-F are scalable for wide datapaths. The throughput and latency behaviour of the SWAG engine can be optimised further. When a higher throughput is needed for large windows, a wider $P$ can be used, or two sorters can be combined and de-multiplexed using rate converters. A lower latency can be achieved by shortening the linear sorter modules, such as by using lower $k$ values for $\textit{WS}\leq k$, or a higher $P$ for $\textit{WS}>k$.
Other future work includes the automation of synthesising customised aggregation engines, such with a high-level synthesis (HLS) library. It could also be explored as a functional unit inside CPUs to work as dedicated SIMD instructions [38].
@@ VII Conclusions
The increasing memory datapath widths are becoming increasingly challenging to exploit with existing programming techniques and models. This includes aggregation, a naturally parallel problem. A modular and highly-adaptable aggregation pipeline is presented that uses prefix-scan and reverse butterfly topologies to perform a wide range of aggregation tasks with high-throughput. The current design improves the theoretical and synthesis-based circuit complexity for specialised group-by applications such as for counting unique elements per group. It is then adapted to work as a competitive sliding window aggregation (SWAG) engine by incorporating a modified sorter that sorts each incoming buffered window on chip. The sorter includes lightweight algorithmic modifications that append group cardinality information to the incoming tuples to assist with median calculation. The resulting designs are shown to have low hardware overheads, while often supporting a superset of the functionality of alternatives. The SWAG engine exhibits up to hundreds of times of speedup when compared to CPU-only approaches on our embedded evaluation platform, and is able to process multiples of the input rate of the state-of-the-art on datacenter-grade FPGAs. This is achieved with the practical assumption of a single window for all group IDs, which in turn has a wider applicability for SWAG workloads. The highly-optimised on-chip pipeline eliminates the need for off-chip memories such as DRAM, and reduces the latency for relatively complex aggregation tasks.

### v4

@@ VI Discussion and Future work
The selection of operators in the aggregation pipeline can easily incorporate other associative operators such as multiplication, bitwise exclusive-OR etc. This can be achieved by including the corresponding logic inside the $n$ and $n^{\prime}$ entities. The average operation is not explored beyond simulation to avoid shifting the focus to floating point division optimisation. This is without loss of generality, only requiring $P$ dividers, usually consuming a few hundred LUT/FFs according to the use case, and without a perceptible timing overhead [divfor].
Similarly, other types of averaging are possible by combining simultaneous associative operators together at the $n^{\prime}$ stage. As an example, combining multiplication and count with a root circuit, can provide the geometric mean. Another instance is with the selection mechanism inside EnthuseSWAP, which can be generalised to support the percentile operator to select arbitrary percentiles based on the pre-calculated cardinality. Other selection queries could use more specialised criteria. Future applications could exploit the additional information in the engine, such as the order of the sorted data, which has already been used in the distinct count algorithm (step $(c)$ in section II-B) to essentially make it an associative operation.
@@
In the comparison of section V-B, a unique aspect for EnthuseSWAG with groups is the single window, while still providing per group statistics. Per-group window SWAG, although supported, might be less trivial to efficiently exploit the presented architecture, as the sorting advantage might diminish when “recycling” more data on each window advance. Thus, it could be seen as an approximation for when the task involves a higher number of groups and there is a need for multiple windows. In practice, keeping live statistics for all groups may not be necessary [6726885], and a simple filtering step to isolate the group IDs would still dedicate the whole window to the desired group.
Yet, the notion of having a common window for grouped data is widely useful for other SWAG tasks. One such case is for supporting more complex SWAG queries without groups, where the group field can carry the value instead, for operating in the context of the values. This includes the mincount or maxcount operator (i.e. counting the instances of the minimum or maximum value within the window) [tangwongsan2021order], which can be implemented as a group-by query where the count operator is only activated on the first or last group, corresponding the minimum or maximum respectively.
@@
The presented engines can be simplified further according to the desired operator selection. If only associative operators (e.g. sum) are needed, then the requirement for the sorted input is only with respect to the group. For SWAG, this essentially means that the comparison width of the small-scale sorter can be constrained to that of the group to minimise its complexity. For non-SWAG, this would also be reflected in the accompanying sorter, or in software equivalents that provide the sorted stream. If selection operators such as the median are not present, then the cardinality logic (section III-D) is not needed. Hence, the disjoint prefix-scan steps can also be merged back. In this scenario, alternative sorters would also suffice for SWAG. When this holds and the input does not carry group information, then a sorter is not needed at all.
As demonstrated in the study of section IV-F are scalable for wide datapaths. The throughput and latency behaviour of EnthuseSWAG can be optimised further. When a higher throughput is needed for large windows, a wider $P$ can be used, or two sorters can be combined and de-multiplexed using rate converters. A lower latency can be achieved by shortening the linear sorter modules, such as by using lower $k$ values for $\textit{WS}\leq k$, or a higher $P$ for $\textit{WS}>k$.
Other future work includes the automation of synthesising customised aggregation engines, such with a high-level synthesis (HLS) library. It could also be explored as a functional unit inside CPUs to work as dedicated SIMD instructions [cui2023risc].
@@ VII Conclusions
The increasing memory datapath widths are becoming increasingly challenging to exploit with existing programming techniques and models. This includes aggregation, a naturally parallel problem. Enthuse, a modular and highly-adaptable aggregation pipeline is presented that uses prefix-scan and reverse butterfly topologies to perform a wide range of aggregation tasks with high-throughput. This design improves the theoretical and synthesis-based circuit complexity for specialised group-by applications such as for counting unique elements per group. It is then adapted to EnthuseSWAG, to work as a competitive sliding window aggregation engine by incorporating a modified sorter that sorts each incoming buffered window on chip. This includes lightweight algorithmic modifications that append group cardinality information to the incoming tuples to assist with median calculation. The resulting designs have low hardware overheads, while often supporting a superset of the functionality of alternatives. EnthuseSWAG exhibits up to hundreds of times of speedup when compared to CPU-only approaches for the same task on our embedded platform. It is also able to process multiples of the input rate of the state-of-the-art. For SWAG with groups, the competitiveness with prior FPGA work is achieved under the practical assumption of a single window. This is already implied for tasks such as SWAG without groups, where EnthuseSWAG processes multiple tuples per cycle without multiple workers. The highly-optimised on-chip pipelines eliminate the need for off-chip memories, and reduce the latency for relatively advanced aggregation tasks.
@@ Acknowledgements
[Uncaptioned image] Philippos Papaphilippou is a Lecturer in Electrical and Electronic Engineering at the University of Southampton. His doctoral studies at Imperial College London (2021) were funded by dunnhumby/Tesco to research novel accelerators for customer analytics. Other industrial engagements included a 3-month contractorship with Microsoft, and a 2-year senior CPU architect position at Huawei. His research topics include parallel algorithms, multi-processors, FPGAs and data science.
[Uncaptioned image] Wayne Luk is a Professor of Computer Engineering at Imperial College London. He leads the Programming Languages and Systems Section, and the Custom Computing Research Group at the Department of Computing. He is a Fellow of the Royal Academy of Engineering, the IEEE, and the BCS.